A particle is projected at an angle of t to the horizontal, with a velocity of u. if the maximum distant travelled is 20m and the greatest height reached is 10m. Find u and t.


1
Expert's answer
2021-01-31T14:03:44-0500

"u_{0x}=u\cdot\cos t"


"u_{0y}=u\cdot\sin t"


"h=u_{0y}t-gt^2\/2"


"u_y=u_{0y}-gt\to t=u_{0y}/g"


"h_{max}=u_{0y}t-gt^2\/2=u_{0y}^2\/g-u_{0y}^2\/(2g)=u_{0y}^2\/(2g)="


"=u^2\cdot\sin^2 t/(2g)"


The maximum distant travelled


"l_{max}=u^2\cdot\sin2t/g" . We get


"h_{max}/l_{max}=\tan t/4\to 10/20=\tan t/4\to \tan t=2"


"t=63.4\u00b0" . Answer


"u=\sqrt{gl_{max}/\sin2t}=\sqrt{9.8\cdot 20/\sin(2\cdot 63.4°)}=15.6(m/s)" . Answer